### v1

@@ 4 Numerical experiments
Table 3 presents results for IC-GMRES-IR using a fp16-$IC(2)$ preconditioner. Results are for $IC(2)$ with look-ahead and the GMW($\beta$) strategy for $\beta=0.5$, 10 and 100. Problem B3 only occurs for GMW(100) (there is a single B3 problem for examples HB/bcsstk24 and Boeing/bcsstk38 and for these a global shift is used). Problem B4 happens only for GMW(10) applied to GHS_psdef/hood (5 occurrences for this example) and GMW(100) applied to HB/bcsstk11 (happens once); again a global shift is used to avoid breakdown. We see that with $\beta=0.5$, for some examples a large number of local modifications ($nmod$) are made. This leads to the preconditioner being of poorer quality compared to the $IC(2)$ preconditioner with look-ahead. For $\beta=10$, local modifications are only needed for a few problems (HB/bcsstk11, HB/bcsstk24 and Um/offshore). In each case, the resulting preconditioner is not successful. For GMW(100), local modifications are only made for Um/offshore; for all other test examples, GMW(100) is equivalent to $IC(2)$ with no look-ahead.
Table 3: Results for IC-GMRES-IR using a fp16-$IC(2)$ preconditioner. Results are for $IC(2)$ with the GMW($\beta$) strategy and $IC(2)$ with look-ahead. $its$ is the number of GMRES iterations and $(n1,nmod)$ are the numbers of times problem B1 is detected and the number of local modifications made by the GMW strategy. For GMW(100), $nmod=0$ in all cases and so is omitted. ${}^{\sharp}$ and ${}^{\ast}$ indicate problems B3 and B4, respectively. $>1000$ indicates the requested accuracy was not obtained within the GMRES iteration limit.
@@
GHS__\__psdef/hood 2459 (0, 25074) ${}^{\ast}$ (0, 0) 581 (5, 0) 581 (5)
Um/offshore $>1000$ (0, 3846) $>1000$ (0, 5838) 2013 (4, 2) 129 (5)
The sensitivity of the GMW($\beta$) approach to the choice of $\beta$ is reported on in Table 4 for problem Boeing/bcsstk38. As $\beta$ increases, the number of local modifications to diagonal entries ($nmod$) steadily decreases and so too does the GMRES iteration count ($its$). For this example, for each $\beta\geq 0.4$, problem B1 was detected once and a global shift $\alpha=10^{-3}$ was then employed.
Table 4: Results for problem Boeing/bcsstk38. IC-GMRES-IR is run using a fp16-$IC(2)$ preconditioner computed with the GMW($\beta$) strategy for a range of values of $\beta$. $nmod$ and $its$ are the numbers of local modifications made by the GMW strategy and GMRES iterations, respectively.
@@
Table 5: Results for IC-GMRES-IR using a fp64-$IC(2)$ preconditioner. Results are for $IC(2)$ with the GMW($\beta$) strategy and $IC(2)$ with look-ahead. $its$ is the number of GMRES iterations and $(n1,nmod)$ are the numbers of times problem B1 is detected and the number of local modifications made by the GMW strategy. For GMW(0.5) and GMW(10), $n1=0$ for all examples so is omitted. $>1000$ indicates the requested accuracy was not obtained within the GMRES iteration limit. ‡ indicates failure to compute useful factors because of enormous growth in the entries.
@@
GHS__\__psdef/hood 2480 (25054) $>1000$ (2013) $>1000$ (0, 11998) 568 (5)
Um/offshore $>1000$ (4094) $>1000$ (6327) ‡(4, 4730) 128 (5)
@@ 5 Concluding remarks
Following on from our earlier study [6], in this paper we have again illustrated the potential for using half precision arithmetic to compute incomplete factorization preconditioners that can be used to recovery double precision accuracy in the solution of highly ill conditioned symmetric positive definite linear systems. In fp16 arithmetic, the danger of breakdown during the factorization of a sparse matrix is imminent and we must employ strategies that force computational robustness. To avoid disastrous breakdown, we have looked at global strategies plus a local modification scheme based on the GMW approach that has been used for dense matrices within the field of optimization. This strategy employs a parameter $\beta$. Choosing a small $\beta$ prevents breakdown during the factorization (in both fp16 and fp64 arithmetic) and there is no need to employ a global shift. However, the penalty for avoiding the need to restart the factorization when potential breakdown is detected is a preconditioner that is of poorer quality than that which is obtained by employing a simple global shifting approach.
When using higher precision arithmetic, the potential dangers within an incomplete factorization algorithm can be hidden. As our experiments have demonstrated, a standard $IC$ factorization using fp64 without look-ahead can lead to an ineffective preconditioner because of growth in the size of the entries in the factors. Without careful monitoring (which is not routinely done), this growth may be unobserved but when subsequently applying the preconditioner, the triangular solves can overflow, resulting in failure (without warning).
Finally, we reiterate that, although our focus has been on symmetric positive definite systems, breakdown and/or large growth in factor entries is also an issue for the incomplete factorization of general sparse matrices. Again, safe checks need to be built into the algorithms and their implementations to guarantee robustness.

### v4

@@ 4 Numerical experiments
Table 3 presents results for IC-GMRES-IR using a fp16-$IC(2)$ preconditioner with look-ahead and the GMW($\beta$) strategy for $\beta=0.5$, 10, and 100. B3 breakdown only occurs for GMW(100) (there is a single B3 breakdown for examples HB/bcsstk24 and Boeing/bcsstk38 and for these a global shift is used). B4 breakdown happens only for GMW(10) applied to GHS_psdef/hood (5 occurrences for this example) and GMW(100) applied to HB/bcsstk11 (happens once); again a global shift is used to avoid breakdown. We see that with $\beta=0.5$, for some examples a large number of local modifications ($nmod$) are made. This leads to the preconditioner being of poorer quality compared to the $IC(2)$ preconditioner with look-ahead. For $\beta=10$, local modifications are only needed for a few problems (HB/bcsstk11, HB/bcsstk24 and Um/offshore). In each case, the resulting preconditioner is not successful. For GMW(100), local modifications are only made for Um/offshore; for all other test examples, GMW(100) is equivalent to $IC(2)$ with no look-ahead.
Table 3: Results for IC-GMRES-IR (Algorithm 5) using a fp16-$IC(2)$ preconditioner with the GMW($\beta$) strategy (Section 3.4) and with look-ahead (Section 3.2). $its$ is the number of GMRES iterations and $(n1,nmod)$ are the numbers of times B1 breakdown is detected and the number of local modifications made by the GMW strategy. and indicate B3 and B4 breakdowns, respectively. NC indicates that on an inner iteration the requested GMRES accuracy of $u_{64}^{1/4}$ was not achieved within the limit of 1000 iterations.
@@
GHS__\__psdef/hood 2459 (0, 25074) 581 (0, 0) 581 (5, 0) 581 (5)
Um/offshore NC (0, 3846) NC (0, 5838) 2013 (4, 2) 129 (5)
The sensitivity of the GMW($\beta$) approach to the choice of $\beta$ is reported on in Table 4 for problem Boeing/bcsstk38. As $\beta$ increases, the number of local modifications to diagonal entries ($nmod$) steadily decreases and so too does the GMRES iteration count ($its$). For this example, for each $\beta\geq 0.4$, B1 breakdown was detected once and a global shift $\alpha=10^{-3}$ was then employed.
Table 4: Results for problem Boeing/bcsstk38. IC-GMRES-IR is run using a fp16-$IC(2)$ preconditioner computed with the GMW($\beta$) strategy for a range of values of $\beta$ (Section 3.4). $nmod$ and $its$ are the numbers of local modifications made by the GMW strategy and GMRES iterations, respectively.
@@
Table 5: Results for IC-GMRES-IR (Algorithm 5) using a fp64-$IC(2)$ preconditioner with the GMW($\beta$) strategy (Section 3.4) and with look-ahead (Section 3.2). $its$ is the number of GMRES iterations and $(n1,nmod)$ are the numbers of times B1 breakdown is detected and the number of local modifications made by the GMW strategy. For GMW(0.5) and GMW(10), $n1=0$ for all examples so is omitted. NC indicates that on an inner iteration the requested GMRES accuracy of $u_{64}^{1/4}$ was not achieved within the limit of 1000 iterations. ‡ indicates failure to compute useful factors because of enormous growth in the entries.
@@
GHS__\__psdef/hood 2480 (25054) NC (2013) NC (0, 11998) 568 (5)
Um/offshore NC (4094) NC (6327) ‡(4, 4730) 128 (5)
All the reported results employed our explicit safe tests for breakdown. We have also run the fp16 arithmetic experiments with the B1 to B4 breakdown tests replaced by IEEE exception handling. As expected, because in fp16 arithmetic $\tau_{u}=10^{-5}$ and $x_{min}=\mathcal{O}(10^{-5})$, this led to the same number of restarts and hence the same iteration counts. However, by only testing the exception flag at the end of each major step of the factorization, this approach did not distinguish between the different types of breakdown. For the experiments using fp64 arithmetic, IEEE exception handling did not detect any problems and consequently, for some examples, this led to growth in the factor entries (exactly as in the case of no look-ahead).
@@ 5 Concluding remarks
Following on from our earlier study [6], in this paper we have illustrated the potential for using half precision arithmetic to compute incomplete factorization preconditioners that can be used to obtain double precision accuracy in the solution of highly ill-conditioned symmetric positive definite linear systems. In fp16 arithmetic, the danger of breakdown during the factorization of a sparse matrix is imminent and we must employ strategies that force computational robustness. To avoid breakdown, we have looked at global strategies plus a local modification scheme based on the GMW approach that has been used for dense matrices within the field of optimization. This employs a parameter $\beta$. Choosing a small $\beta$ prevents breakdown during the factorization (in both fp16 and fp64 arithmetic) and there is no need to employ a global shift. However, the penalty is of poorer quality than that which is obtained by employing a simple global shifting approach. Thus, our recommendations are to always prescale the problem, to use a global shift, and to incorporate look-ahead. In addition, when developing software using fp16, monitoring for breakdown must be built in to ensure robustness. If this is done, then using low precision to compute an effective preconditioner appears to be feasible.
Once a Fortran compiler that supports bfloat16 becomes available, it would be very interesting to compare its performance to that of fp16. bfloat16 has the same exponent size as fp32 (single precision). Consequently, converting from fp32 to bfloat16 is easy: the exponent is kept the same and the significand is rounded or truncated from 24 bits to 8; hence overflow and underflow are not possible in the conversion. The disadvantage of bfloat16 is its lesser precision: essentially 3 significant decimal digits versus 4 for fp16. Another possible future direction is to explore the effects of different preorderings of $A$ on the number of breakdowns and the quality of the low precision factors. Fill-reducing orderings can result in later entries in the factor being updated by more entries from the previous columns. Intuitively, this may lead to more breakdowns.
When using higher precision arithmetic, the potential dangers within an incomplete factorization algorithm can be hidden. As our experiments have demonstrated, a standard $IC$ factorization using fp64 arithmetic without look-ahead can lead to an ineffective preconditioner because of growth in the size of the entries in the factors. Without careful monitoring (which is not routinely done), this growth may be unobserved but when subsequently applying the preconditioner, the triangular solves can overflow, resulting in the computation aborting.
Finally, we reiterate that, although our focus has been on symmetric positive definite systems, breakdown and/or large growth in factor entries is also an issue for the incomplete factorization of general sparse matrices. Again, safe checks (or the use of IEEE exception handling) need to be built into the algorithms and their implementations to guarantee robustness.
[…]
Acknowledgements
We would like to thank John Reid for discussions on exception handling in Fortran and for carefully reading and commenting on a draft of this paper. We are also grateful to two anonymous referees for their insightful feedback that has led to improvements in the paper.
Declarations
  • Funding. The authors received no external grant funding for this research.
  • Ethics approval and consent to participate. There was no ethics approval required for this research.
  • Data. The test matrices used in this study are taken from the SuiteSparse Collection and are available at https://sparse.tamu.edu/